# Physics Context Builders: A Modular Framework for Physical Reasoning in Vision-Language Models

### v2

@@ A.2 PCB training
    Human-Narration (HN), which generates a summary for all the collisions happened in the scene.
@@
    Structured-Physics (SP), which describes each provided video frame separately as follows, while adding physical properties of the objects, including their discretized and normalized locations and velocities. We also include the location of the collisions happened up to a certain frame.
@@
Training details. We use the pre-trained PaliGemma-3B model for training the PCB modules and apply the LoRA fine-tuning scheme, similar to the approach used in the QA fine-tuning. Table 7 (right) provides the hyper-parameters used to train PCB modules for both Falling Tower and CLEVRER datasets.
@@ B.1 The Effect of Framing Multi-Choice Questions as Multiple Binary Questions
As discussed in the main paper, framing the multi-choice questions as multiple binary questions in CLEVRER can yield significant improvement in the accuracy of the models. In Table 8, we provide a comparison between the performance of fine-tuned PaliGemma-3B models with and without this change. As demonstrated, we observe improvement in almost all categories, except for the per question predictive accuracy. We posit that this is because the predictive questions in CLEVRER are always binary questions with exactly one correct choice. Framing the predictive questions as two independent binary questions can result in a model choosing both options as correct or wrong.
@@ B.2 The Effect of Training Epochs
We run an ablation study to assess the effect of training for smaller vs. larger number of epochs on the accuracy of CLEVRER in the QA fine-tuning task. Table 9 demonstrates a large improvement in training for more epochs.
Table 9: The performance of the fine-tuned PaliGemma-3B model on question answer pairs for the CLEVRER benchmark based on the number of trained epochs. Here, multi-choise questions are asked as they are (without framing them as multiple binary questions).
@@ B.3 Evaluating the Importance of Vision Module
We illustrate the importance of the vision module in a VLM for physical reasoning by conducting the following experiment. Here, we QA-fine-tune only the language model part of PaliGemma-3B while freezing the vision module. The results in Table 10 shows that the performance across all categories drops slightly for the language model-only setting. Therefore, jointly fine-tuning both the vision and language modules is essential for optimal performance, as it enables the model to better align visual features with linguistic representations.
Table 10: Performance drop due to freezing the vision module on the PaliGemma3b-base model for the QA fine-tuning over CLEVRER.
@@ Appendix C Falling Tower Dataset
The Falling Tower dataset is a benchmark for stability detection of stacked objects, inspired by the ShapeStacks benchmark Groth et al., (2018). It includes 4864 unique scenes, 72,775 questions, and detailed simulation-generated annotations to support training Vision-Language Models (VLMs) for spatial and physical reasoning. Each simulation instance is represented as a JSON file containing:
@@
Figure 6 shows the distribution of object stacks in terms of their stability, as well as the distribution of question types.

### v3

@@ A.2 PCB training
    Human-Narration (HN), which generates a summary of all the collisions that occurred in the scene.
@@
    Structured-Physics (SP), which describes each provided video frame separately as follows, while adding physical properties of the objects, including their discretized and normalized locations and velocities. We also include the locations of collisions that occurred up to a certain frame.
@@
Training details. We use the pre-trained PaliGemma-3B model for training the PCB modules and apply the LoRA fine-tuning scheme, similar to the approach used for QA fine-tuning. Tab. 7 (right) provides the hyperparameters used to train PCB modules for both Falling Tower and CLEVRER datasets.
@@ B.1 The Effect of Framing Multi-Choice Questions as Multiple Binary Questions
As discussed in the main paper, framing the multi-choice questions as multiple binary questions in CLEVRER can yield significant improvement in the accuracy of the models. In Tab. 8, we provide a comparison between the performance of fine-tuned PaliGemma-3B models with and without this change. As demonstrated, we observe improvement in almost all categories, except for the per question predictive accuracy. We posit that this is because the predictive questions in CLEVRER are always binary questions with exactly one correct choice. Framing the predictive questions as two independent binary questions can result in a model choosing both options as correct or wrong.
@@ B.2 The Effect of Training Epochs
We run an ablation study to assess the effect of training for smaller vs. larger number of epochs on the accuracy of CLEVRER in the QA fine-tuning task. Tab. 9 demonstrates a large improvement in training for more epochs.
Table 9: The performance of the fine-tuned PaliGemma-3B model on question answer pairs for the CLEVRER benchmark based on the number of trained epochs. Here, multi-choice questions are asked as they are (without framing them as multiple binary questions).
@@ B.3 Evaluating the Importance of Vision Module
We illustrate the importance of the vision module in a VLM for physical reasoning by conducting the following experiment. Here, we QA-fine-tune only the language model part of PaliGemma-3B while freezing the vision module. The results in Tab. 10 shows that the performance across all categories drops slightly for the language model-only setting. Therefore, jointly fine-tuning both the vision and language modules is essential for optimal performance, as it enables the model to better align visual features with linguistic representations.
Table 10: Performance drop due to freezing the vision module on the PaliGemma-3B-base model for the QA fine-tuning over CLEVRER.
@@ Appendix C Falling Tower Dataset
The Falling Tower dataset is a benchmark for stability detection of stacked objects, inspired by the ShapeStacks benchmark [15]. It includes 4864 unique scenes, 72,775 questions, and detailed simulation-generated annotations to support training Vision-Language Models (VLMs) for spatial and physical reasoning. Each simulation instance is represented as a JSON file containing:
@@
Fig. 5 shows the distribution of object stacks in terms of their stability, as well as the distribution of question types.
@@
Appendix D Additional Experiments
D.1 Specialized Baselines for CLEVRER
Here, we compare the fine-tuned PaliGemma-3B model on the CLEVRER QA dataset to specialized architectures designed specifically for CLEVRER. Although the fine-tuned model does not outperform all benchmarks, its comparable performance highlights the potential benefits of generalist models over bespoke baselines.
Table 11: Per-question performance of fine-tuned PaliGemma-3B compared to specialized methods on CLEVRER.
Category Model Descriptive Explanatory Predictive Counterfactual Specialized Methods VRDP [12] 89.80 82.40 83.80 75.70 DCL [9] 90.70 82.80 82.00 46.50 CRCG [20] 95.55 99.81 76.64 78.31 Aloe [11] 94.00 96.00 87.50 75.60 Fine-tuned QA PaliGemma-3B-Fine-Tuned 92.90 94.70 83.60 68.40
D.2 The Effect of PCBs on the InternVL 3.0 Model
Table 12: Performance of InternVL 3.0 (8B parameters), augmented with Physics Context Builders (PCBs), compared to its zero-shot version on the Falling Tower benchmark. HN refers to the Human Narration-style PCB. The second value after the slash indicates the Sim2Real accuracy, and the third value represents the F1 score on Sim2Real.
Model Descriptive [sim acc. / real acc.] Stability [sim acc. / real acc. / real F1] num. obj. num. obj. $\uparrow\downarrow$ obj. $\uparrow\downarrow$ obj. stable tower stable InternVL3-8B 81.57 / 78.95 52.77 / 78.95 53.85 / 84.21 52.71 / 84.21 / 80.19 46.42 / 73.68 / 68.64 InternVL3-8B-PCB 95.94 / 88.24 66.29 / 70.58 70.01 / 94.12 69.21 / 76.47 / 73.39 83.07 / 76.47 / 73.73